OpenStudy (anonymous):

A farmer sells strawberris at a market in both pint-sized and quart-sized containers. The farmer charges $3 for each pint, and $5 for each quart and is always paid on the day of purchase and sells no other goods. On a recent day, the farmer sold as many pint containers as quart containers and recieved $120 in sales. How many pints of strawberries did the farmer sell?

hero (hero):

Hint: x = y 3x + 5y = 120

hero (hero):

I think that first equation I posted was too much of a give away.

hero (hero):

That's really all you need to know. If x = y, then just use substitution.

OpenStudy (anonymous):

The number of pints he sells = p The number of quarts he sells = q at $3 per pint, the total money he will make on pints are $3*p = $3p at $5 per quart, the total money he will make on quarts are $5*q = $5q The total he got that day was 120 meaning that $3p + $5q = 120 but since he sold an equal number of pints as quarts then p = q that day so that day, $3p + $5p = 120 $8p = 120 p = 15